OpenStudy (anonymous):

The length of a rectangle is five feet longer than its width. If the area is 84 square feet, find the side lengths.

OpenStudy (shamil98):

lw = 84 l = 5 + w w = l - 5 w(5+w) = 84 w^2 + 5w - 84 = 0 use the quadratic formula.